# Bayesian Optimization with Gaussian Processes to Accelerate Stationary Point Searches

### 16 Mar 2026 v2

@@ 3.1  What the Surrogate Must Provide
The surrogate model must interpolate a small, incrementally growing set of energy and force data, provide calibrated uncertainty for active learning, and remain cheap enough that fitting and querying costs far less than the electronic structure call it replaces. Gaussian process regression satisfies all three. This section develops the GP from the perspective of building and using the surrogate, covering what quantities need to be computed, how they connect to the physics, and where the bottlenecks arise. We refer readers to Rasmussen and Williams [81] and Gramacy [35] for the mathematical foundations. As noted earlier, Deringer, Bartok, and Csanyi [17] has a detailed review of GPR in atomistic simulation from a global MLIP view, including structural descriptors (SOAP, ACE), sparse approximations, and validation methodology. That review treats the GP as a tool for building global machine-learned potentials from large databases; the present treatment focuses on the complementary regime of local surrogates built on the fly from tens of data points. The key distinction is hyperparameter management: MLIP approaches optimize hyperparameters once on a large training set and fix them, while the local GP re-optimizes at every step as the training set grows, requiring trust regions and active data selection to maintain stability.
@@
The posterior mean (Eq. 3.3) coincides with the kernel ridge regression (KRR) estimator with regularization $\sigma_{n}^{2}$ [81], but the GP additionally provides the predictive variance (Eq. 3.4), which is the basis for the acquisition criterion in Section 6.
@@
The interaction between these two quantities during a search is illustrated by the following progression. In the first few iterations, the GP has little data and the variance is large everywhere except at the evaluated configurations (Figure 3, top-left panel). The surrogate prediction is correspondingly uncertain, and the trust region (Section 5.2) constrains the optimizer to small steps. As data accumulates, the variance shrinks in the neighborhood of the reaction path (Figure 4) and the surrogate becomes a faithful replica of the true PES in that local region (Figure 3, bottom panels). The optimizer can now take longer steps on the cheap surrogate, and the active learning criterion directs the next expensive evaluation to the frontier where the variance is still large. This feedback between uncertainty, data acquisition, and optimization step length accounts for the roughly ten-fold savings in electronic structure calls.
Figure 3: GP surrogate fidelity as a function of training set size on the Muller-Brown surface. Each panel shows the GP posterior mean contours after training on $M=5,15,30,50$ Latin hypercube-sampled configurations (black dots). With 5 points the surrogate captures only crude basin structure; by 30 points the contours closely match the true PES (Figure 12) in the sampled region.
Figure 4: GP predictive variance on the Muller-Brown surface after 20 training evaluations clustered near minimum A and saddle S1 (black dots). The variance is near zero close to training data and grows with distance, reaching a maximum (coral diamond) in the unexplored region. This variance landscape is the basis for active learning: the next electronic structure evaluation is placed where the GP is least certain.
@@ 3.3.1 The Inverse-Distance Squared Exponential Kernel
The length-scale parameters $l_{\phi(i,j)}$ control the GP’s sensitivity to changes in each interatomic distance. A short length scale for a particular atom pair means the GP treats small changes in that pair’s inverse distance as significant (i.e., the pair is “stiff” in the model’s view); a long length scale means the GP is insensitive to that pair. In practice, the hyperparameter optimization (Section 3.4) learns these from the data, and bonds that are actively breaking or forming during the reaction acquire short length scales, while spectator bonds that barely change acquire long ones. This automatic relevance determination is what allows the GP to focus its limited training data on the degrees of freedom that matter for the particular transition being studied.
The constant offset $\sigma_{c}^{2}$ is fixed rather than optimized alongside the other hyperparameters, since with the small training sets typical of on-the-fly searches ($M\sim 10\textrm{--}50$), the marginal likelihood cannot reliably distinguish $\sigma_{c}^{2}$ from $\sigma_{f}^{2}$. A default of $\sigma_{c}^{2}=1.0$ works well for molecular systems with eV-scale energies; for 2D model surfaces (LEPS, Muller-Brown) where energies are already centered near zero, $\sigma_{c}^{2}=0.0$ avoids introducing a superfluous degree of freedom. Without it, the GP prior mean is zero, and the posterior mean would revert to zero far from the training data. The constant kernel adds a baseline covariance that is independent of configuration, which allows the GP to represent a nonzero mean energy level. In practice, this absorbs the large absolute energy that is common in electronic structure calculations, so the GP only needs to model the relative energy variations. Because the constant kernel contributes only to the energy-energy block of the covariance matrix (its derivative with respect to any coordinate is zero), it does not affect force predictions and carries no risk of corrupting the gradient information that drives the optimization.
@@ 3.4  Hyperparameter Optimization
The necessity of per-step hyperparameter optimization is a distinguishing feature of the local surrogate regime that sets it apart from global GP potentials. In MLIP frameworks, especially universal models [64] the kernel operates in a descriptor space (SOAP, ACE) whose structure already encodes the relevant chemical environment. The descriptors are designed so that the kernel length scale has a physical interpretation (the Gaussian mollification width $\sigma_{a}$, typically 0.3 Å for systems containing hydrogen and 0.5 Å for heavier elements) that is nearly universal across chemical systems. The practitioner fixes these hyperparameters a priori from physical reasoning and then builds the training set to achieve a target accuracy, rather than optimizing hyperparameters to match a fixed dataset. The regularization parameter is set to the estimated noise level of the training data, and data is added until the model reaches this noise floor. The problem is, as it were, “turned upside down”: instead of fitting hyperparameters to data, one chooses hyperparameters that encode prior physical knowledge and fits the data to the model.
@@
The signal variance $\sigma_{f}^{2}$ can be handled in multiple ways. In this framework it is a free hyperparameter optimized by MLL, with a logarithmic barrier (Section 8.2) to prevent divergence. An alternative is to marginalize $\sigma_{f}^{2}$ out under a conjugate inverse-gamma prior. The result is a Student’s t-process [88] whose predictive mean is identical to the GP’s but whose heavier-tailed variance produces more conservative uncertainty estimates; this has been applied to surrogate-accelerated NEB for surface catalysis [99]. Other kernels (e.g., Matern [15]) can each be made to work with appropriate tuning. The requirements are that the surrogate is locally faithful, that the uncertainty is calibrated well enough to guide sampling, and that the hyperparameters do not destabilize the model; the specific mechanism for achieving these is secondary.
@@ 4 The Bayesian Surrogate Loop: Anatomy of the Unified Framework
Input: Initial configuration(s) $\mathbf{X}_{0}$, oracle $V(\cdot)$, convergence threshold $\epsilon$
1 Evaluate oracle at $\mathbf{X}_{0}$; initialize $\mathcal{D}=\{(\mathbf{x},V(\mathbf{x}),\nabla V(\mathbf{x}))\}$;
2 repeat
3 Select training subset $\mathcal{S}\subseteq\mathcal{D}$;
4 Train hyperparameters $\boldsymbol{\theta}$ on $\mathcal{S}$;
5 Build prediction model from $\mathcal{D}$ with $\boldsymbol{\theta}$;
6 Optimize on surrogate $V_{\text{GP}}$: method-specific inner loop;
7 Clip proposed step via trust region;
8 Acquire: select next oracle point $\mathbf{x}^{*}$ via criterion $\alpha(\mathbf{x})$;
9   Evaluate oracle at $\mathbf{x}^{*}$; $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\mathbf{x}^{*},V,\nabla V)\}$;
10
11until $|\mathbf{F}(\mathbf{x}^{*})|<\epsilon$;
@@ 5.1.2 Phase 2: GPR Iterations (Rotation + Translation on the Surrogate)
Input: Training set $\mathcal{D}$ from Phase 1, force threshold $\epsilon_{\text{force}}$
Output: saddle point at $\mathbf{R}_{0}$
1 repeat
2   Fit GP hyperparameters by maximizing $\mathcal{L}(\boldsymbol{\theta})$ (Eq. 3.14) on $\mathcal{D}$;
3   Run full dimer (rotation via CG + translation via L-BFGS) on $V_{\text{GP}}(\mathbf{x})$;
 // until surrogate convergence or trust radius violated
4   Evaluate true PES at midpoint $\mathbf{R}_{0}$;
5 if $|\mathbf{F}(\mathbf{R}_{0})|<\epsilon_{\text{force}}$ then
6    return saddle point at $\mathbf{R}_{0}$;
7
8   end if
9  Add $(\mathbf{x},V,\nabla V)$ to $\mathcal{D}$;
10
11until converged;
@@ 9  Add $(\mathbf{x},V,\nabla V)$ to $\mathcal{D}$;
The source of the savings is easy to trace. In the “improved” dimer formulation in eOn [12]  [^2] https://eondocs.org [51, 74], each translation step requires 5 to 15 rotation evaluations (each a full electronic structure call) to converge the orientation. The GP replaces almost all of these inner rotations with surrogate queries that cost microseconds rather than minutes, though the first few initial rotations take true forces. The outer loop still needs one true evaluation per translation step to validate the surrogate and extend the training set, but the inner loop is essentially free. Benchmarks [30, 29] show the median evaluation count dropping by a factor of ten which is precisely the factor one expects from eliminating the inner rotation cost. Despite working in Cartesian coordinates, the GP-dimer achieves performance comparable to internal-coordinate methods (Sella [45]), because the inverse-distance kernel (Eq. 3.10) learns per-pair length scales that adapt to the stiffness landscape of each reaction.
@@ 5.2.2 Atoms Approaching Too Closely
Figure 11: Trust region mechanism on a 1D slice ($y=0.5$) of the Muller-Brown surface. The GP posterior mean (teal) and $\pm 2\sigma$ confidence band (light blue) are accurate near the training data (black dots) but diverge from the true surface (black dashed) outside the trust boundaries (magenta dotted verticals). A hypothetical GP-proposed step at $x=1.0$ (coral cross, labeled “GP step”) falls outside the trust region, where the surrogate is unreliable; the algorithm instead evaluates the true PES at the trust boundary (teal star, “Oracle fallback”).

### 16 Apr 2026 v3

@@ 3.1  What the Surrogate Must Provide
The surrogate model must interpolate a small, incrementally growing set of energy and force data, provide a posterior-variance signal that shrinks where the model has seen data and grows elsewhere (noting that this signal measures sampling density in the kernel geometry, not accuracy against the true PES), and remain cheap enough that fitting and querying costs far less than the electronic structure call it replaces. Gaussian process regression satisfies all three. This section develops the GP from the perspective of building and using the surrogate, covering what quantities need to be computed, how they connect to the physics, and where the bottlenecks arise. We refer readers to Rasmussen and Williams [83] and Gramacy [36] for the mathematical foundations. As noted earlier, Deringer, Bartok, and Csanyi [18] has a detailed review of GPR in atomistic simulation from a global MLIP view, including structural descriptors (SOAP, ACE), sparse approximations, and validation methodology. That review treats the GP as a tool for building global machine-learned potentials from large databases; the present treatment focuses on the complementary regime of local surrogates built on the fly from tens of data points. The key distinction is hyperparameter management: MLIP approaches optimize hyperparameters once on a large training set and fix them, while the local GP re-optimizes at every step as the training set grows, requiring trust regions and active data selection to maintain stability.
@@
The posterior mean (Eq. 3.3) coincides with the kernel ridge regression (KRR) estimator with regularization $\sigma_{n}^{2}$ [83], but the GP additionally provides the predictive variance (Eq. 3.4), which is the basis for the acquisition criterion in Section 6. Before proceeding, it is worth stating clearly what the predictive variance is and what it is not, because the distinction governs how every acquisition rule in this review should be read. Inspecting Eq. 3.4, $\sigma^{2}(\mathbf{x}_{*})=k(\mathbf{x}_{*},\mathbf{x}_{*})-\mathbf{k}_{*}^{\top}\mathbf{K}^{-1}\mathbf{k}_{*}$, the formula depends only on the kernel, the training locations, and the hyperparameters; the observed energies $\mathbf{y}$ enter only through the posterior mean, not through the variance. By construction, $\sigma^{2}$ collapses to the noise floor at every training point and shrinks monotonically in a kernel-dependent neighbourhood of those points as the dataset grows. A low $\sigma^{2}$ therefore reports that $\mathbf{x}_{*}$ is close to the training set in the geometry that the kernel defines; it does not compare the posterior mean against the true PES and is not a calibrated measure of accuracy.
The spline analogy due to Wahba [103] sharpens the point. A GP with kernel $k$ and noise $\sigma_{n}^{2}$ has the same MAP predictor as a smoothing spline that minimizes $\sum_{i}(y_{i}-f(\mathbf{x}_{i}))^{2}/\sigma_{n}^{2}+\|f\|_{\mathcal{H}_{k}}^{2}$, where $\|\cdot\|_{\mathcal{H}_{k}}$ is the reproducing-kernel Hilbert-space norm induced by $k$. Viewed through that lens, $\sigma^{2}(\mathbf{x}_{*})$ is a kernel-space interpolation radius around $\mathbf{x}_{*}$: the quantity a spline-theorist would write down to measure how close $\mathbf{x}_{*}$ is to the scattered data locations $\{\mathbf{x}_{i}\}$. It is exactly as informative as the following heuristic: pass a cubic spline through a cluster of three closely spaced points and a second cluster three Angstroms away; the spline self-report of "error" is small everywhere inside each cluster and between consecutive knots, but the fit to a function that does not actually live in the spline’s smoothness class can be arbitrarily bad in the gaps. The GP inherits the same limitation. Accuracy against the true PES is the unknown the search is trying to resolve; it is simply not part of the quantities the GP computes until the oracle is called.
For a per-search GP this gap is sharper still, because the kernel length scales and signal variance are fit from the same handful of observations whose variance we then compute. The resulting signal is self-referential: sampling more drives $\sigma^{2}$ down by construction, independently of whether the mean has converged to the truth. The active learning criteria built on $\sigma^{2}$ in Sections 5 and 6 are therefore best read as sampling-density signals that complement, rather than replace, geometric safeguards like the trust region.
@@
The interaction between these two quantities during a search is illustrated by the following progression. In the first few iterations, the GP has little data and the variance is large everywhere except at the evaluated configurations (Figure 3, top-left panel). The surrogate prediction is correspondingly uncertain, and the trust region (Section 5.2) constrains the optimizer to small steps. As data accumulates, the variance shrinks in the neighborhood of the reaction path (Figure 4) and the surrogate becomes a faithful replica of the true PES in that local region (Figure 3, bottom panels). The optimizer can now take longer steps on the cheap surrogate, and the active learning criterion directs the next expensive evaluation to the frontier where the variance is still large. This feedback between uncertainty, data acquisition, and optimization step length accounts for the roughly ten-fold savings in electronic structure calls reported on saddle searches in the production papers [31, 30]. These factor-of-ten gains depend on three conditions: an oracle that dominates the per-call cost (so amortized GP overhead remains negligible), an initial configuration far enough from the target that the inner surrogate-driven steps replace many true-PES steps, and access to analytical forces (which provide $3N+1$ data points per call rather than one). Minimization near a quadratic basin gains less, because L-BFGS already converges in few steps; saddle searches with steep, anisotropic regions gain the most. Section 7 revisits this trade-off quantitatively for the LEPS surface.
Figure 3: GP surrogate fidelity as a function of training set size on the Muller-Brown surface. Each panel shows the GP posterior mean contours after training on $N=3,8,15,30$ Latin hypercube-sampled configurations (white markers). With three points the surrogate captures only crude basin structure; by 30 points the contours closely match the true PES (Figure 12) in the sampled region.
Figure 4: GP predictive variance on the Muller-Brown surface after 20 training evaluations clustered near minimum A and saddle S1 (black dots). The variance is near zero close to training data and grows with distance, reaching a maximum (coral diamond) in the unexplored region. This landscape is a map of sampling density in the kernel geometry, not of accuracy against the true surface: it tells the active-learning loop where the surrogate has seen least data, which is the quantity an acquisition rule needs, but a low-variance prediction does not by itself imply a small error.
@@ 3.3  Covariance Functions for Molecular Systems
The inverse-distance feature map remains well-defined for any geometry without coincident atoms, including planar molecules and linear arrangements. The Jacobian
$\frac{\partial\phi_{ij}}{\partial z^{(k)}}\;=\;-\,\frac{z^{(i)}-z^{(j)}}{r_{ij}^{\,3}}\,(\delta_{ki}-\delta_{kj})$ (3.9)
vanishes identically for every $(i,j,k)$ when every atom satisfies $z^{(k)}=0$, and by the chain rule the GP out-of-plane force $F_{z}^{(k)}=-\partial V_{\text{GP}}/\partial z^{(k)}$ is zero at the planar geometry independently of the GP coefficients. This is the physically correct answer rather than a defect: a perpendicular displacement of magnitude $\delta z$ gives $r_{ij}(\delta z)=r_{ij}(0)+\mathcal{O}(\delta z^{2})$, so the energy is genuinely stationary in the symmetry direction, and any function that depends on $\mathbf{x}$ only through $\{r_{ij}\}$ has identically zero gradient along the planar orbit. The instant any atom is perturbed off the plane $z^{(i)}-z^{(j)}$ is generically nonzero, the out-of-plane Jacobian block recovers full rank, and the GP regains sensitivity to all three Cartesian directions.
@@ 3.3.1 The Inverse-Distance Squared Exponential Kernel
The length-scale parameters $l_{\phi(i,j)}$ control the GP’s sensitivity to changes in each interatomic distance. A short length scale for a particular atom pair means the GP treats small changes in that pair’s inverse distance as significant (i.e., the pair is "stiff" in the model’s view); a long length scale means the GP is insensitive to that pair. In practice, the hyperparameter optimization (Section 3.4) learns these from the data, and bonds that are actively breaking or forming during the reaction acquire short length scales, while spectator bonds that barely change acquire long ones. This automatic relevance determination is what allows the GP to focus its limited training data on the degrees of freedom that matter for the particular transition being studied.
The constant offset $\sigma_{c}^{2}$ is fixed rather than optimized alongside the other hyperparameters, since with the small training sets typical of on-the-fly searches ($M\sim 10\textrm{--}50$), the marginal likelihood cannot reliably distinguish $\sigma_{c}^{2}$ from $\sigma_{f}^{2}$. A default of $\sigma_{c}^{2}=1.0$ works well for molecular systems with eV-scale energies; for 2D model surfaces (LEPS, Muller-Brown) where energies are already centered near zero, $\sigma_{c}^{2}=0.0$ avoids introducing a superfluous degree of freedom. Without it, the GP prior mean is zero, and the posterior mean would revert to zero far from the training data. The constant kernel adds a baseline covariance that is independent of configuration, which allows the GP to represent a nonzero mean energy level. In practice this absorbs the large absolute energy common in electronic structure calculations, so the GP only needs to model the relative energy variations. The constant kernel carries zero derivative with respect to any coordinate, so $\sigma_{c}^{2}$ drops out of every derivative block of the covariance matrix and out of the cross-covariance vector $\mathbf{k}_{*}$ used for force prediction; the analytical force expression contains no explicit $\sigma_{c}^{2}$ term. The decoupling stops there. The GP weights come from a single joint solve over the full covariance matrix, so changing $\sigma_{c}^{2}$ shifts the conditioning of that solve and rebalances the energy and force residuals it minimizes; an inappropriate value can still perturb the predicted forces indirectly. The defaults above keep this indirect effect small in practice.
@@ 3.4  Hyperparameter Optimization
The necessity of per-step hyperparameter optimization is a distinguishing feature of the local surrogate regime that sets it apart from global GP potentials. In MLIP frameworks, especially universal models [65] the kernel operates in a descriptor space (SOAP, ACE) whose structure already encodes the relevant chemical environment. The descriptors are designed so that the kernel length scale has a physical interpretation (the Gaussian mollification width $\sigma_{a}$, typically 0.3 Å for systems containing hydrogen and 0.5 Å for heavier elements) that is nearly universal across chemical systems. The practitioner fixes these hyperparameters a priori from physical reasoning and then builds the training set to achieve a target accuracy, rather than optimizing hyperparameters to match a fixed dataset. The regularization parameter is set to the estimated noise level of the training data, and data is added until the model reaches this noise floor. The problem is, as it were, "turned upside down": instead of fitting hyperparameters to data, one chooses hyperparameters that encode prior physical knowledge and fits the data to the model.
@@
The signal variance $\sigma_{f}^{2}$ can be handled in multiple ways. In this framework it is a free hyperparameter optimized by MLL, with a logarithmic barrier (Section 8.2) to prevent divergence. An alternative is to marginalize $\sigma_{f}^{2}$ out under a conjugate inverse-gamma prior. The result is a Student’s t-process [90] whose predictive mean is identical to the GP’s but whose heavier-tailed variance produces more conservative uncertainty estimates; this has been applied to surrogate-accelerated NEB for surface catalysis [102]. Other kernels (e.g., Matern [16]) can each be made to work with appropriate tuning. The requirements are that the surrogate is locally faithful, that the posterior variance is a usable sampling-density signal (even though it does not directly measure accuracy), and that the hyperparameters do not destabilize the model; the specific mechanism for achieving these is secondary.
@@ 4 The Bayesian Surrogate Loop: Anatomy of the Unified Framework
Core versus extensions.
The core components used by every method in this review are the SE kernel with inverse-distance features (Section 3), gradient observations in the covariance matrix (Eq. 3.8), farthest-point sampling for training subsets, MAP regularization of hyperparameters, the adaptive trust radius, and the LCB inner-loop convergence criterion (Eq. 4.2). The OT-GP extensions that further improve accuracy and stability for harder problems are per-bead FPS with Earth Mover’s Distance, the variance barrier and oscillation detection, and the OTGPD adaptive inner tolerance. A separate optional extension is the random Fourier feature approximation (Section 8.4). Section 8 develops each subsection by leading with the core formulation and flagging the OT-GP refinement; readers who want the unified loop without OT-GP can follow only the core text in each subsection.
@@
1:Initial configuration(s) $\mathbf{X}_{0}$, oracle $V(\cdot)$, convergence threshold $\epsilon$
2:Evaluate oracle at $\mathbf{X}_{0}$; initialize $\mathcal{D}=\{(\mathbf{x},V(\mathbf{x}),\nabla V(\mathbf{x}))\}$
3:repeat
4:  Select training subset $\mathcal{S}\subseteq\mathcal{D}$ $\triangleright$ FPS, Section 8.1
5:  Train hyperparameters $\boldsymbol{\theta}$ on $\mathcal{S}$ $\triangleright$ SCG on MAP-NLL, Section 8.2
6:  Build prediction model from $\mathcal{D}$ with $\boldsymbol{\theta}$ $\triangleright$ exact GP or RFF, Section 8.4
7:  Optimize on surrogate $V_{\text{GP}}$: method-specific inner loop $\triangleright$ step 5
8:  Clip proposed step via trust region $\triangleright$ Section 8.3
9:  Acquire: select next oracle point $\mathbf{x}^{*}$ via criterion $\alpha(\mathbf{x})$ $\triangleright$ step 7
10:  Evaluate oracle at $\mathbf{x}^{*}$; $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\mathbf{x}^{*},V,\nabla V)\}$
11:until $|\mathbf{F}(\mathbf{x}^{*})|<\epsilon$
@@ 5.1.2 Phase 2: GPR Iterations (Rotation + Translation on the Surrogate)
1:Training set $\mathcal{D}$ from Phase 1, force threshold $\epsilon_{\text{force}}$
2:repeat
3:  Fit GP hyperparameters by maximizing $\mathcal{L}(\boldsymbol{\theta})$ (Eq. 3.15) on $\mathcal{D}$
4:  Run full dimer (rotation via CG + translation via L-BFGS) on $V_{\text{GP}}(\mathbf{x})$
5:   until surrogate convergence or trust radius violated
6:  Evaluate true PES at midpoint $\mathbf{R}_{0}$
7:  if $|\mathbf{F}(\mathbf{R}_{0})|<\epsilon_{\text{force}}$ then
8:   return saddle point at $\mathbf{R}_{0}$
9:  end if
10:  Add $(\mathbf{x},V,\nabla V)$ to $\mathcal{D}$
11:until converged
@@
The source of the savings is easy to trace. In the "improved" dimer formulation in eOn [13]  [^2] https://eondocs.org [52, 75], each translation step requires 5 to 15 rotation evaluations (each a full electronic structure call) to converge the orientation. The GP replaces almost all of these inner rotations with surrogate queries that cost microseconds rather than minutes, though the first few initial rotations take true forces. The outer loop still needs one true evaluation per translation step to validate the surrogate and extend the training set, but the inner loop is essentially free. Benchmarks [31, 30] show the median evaluation count dropping by a factor of ten which is precisely the factor one expects from eliminating the inner rotation cost. Despite working in Cartesian coordinates, the GP-dimer achieves performance comparable to internal-coordinate methods (Sella [46]), because the inverse-distance kernel (Eq. 3.11) learns per-pair length scales that adapt to the stiffness landscape of each reaction.
@@ 5.2.2 Atoms Approaching Too Closely
Figure 11: Trust region mechanism on a 1D slice ($y=0.5$) of the Muller-Brown surface. The GP posterior mean (teal) and $\pm 2\sigma$ confidence band (light blue) are accurate near the training data (black dots) but diverge from the true surface (black dashed) outside the trust boundaries (magenta dotted verticals). A hypothetical GP-proposed step at $x=1.0$ (coral cross, labeled "GP step") falls outside the trust region, where the surrogate is unreliable; the algorithm instead evaluates the true PES at the trust boundary (teal star, "Oracle fallback").